# BEATS: Optimizing LLM Mathematical Capabilities with BackVerify and Adaptive Disambiguate based Efficient Tree Search

### v1

@@ Abstract
Large Language Models (LLMs) have exhibited exceptional performance across a broad range of tasks and domains. However, they still encounter difficulties in solving mathematical problems due to the rigorous and logical nature of mathematics. Previous studies have employed techniques such as supervised fine-tuning (SFT), prompt engineering, and search-based methods to improve the mathematical problem-solving abilities of LLMs. Despite these efforts, their performance remains suboptimal and demands substantial computational resources. To address this issue, we propose a novel approach, BEATS, to enhance mathematical problem-solving abilities. Our method leverages newly designed prompts that guide the model to iteratively rewrite, advance by one step, and generate answers based on previous steps. Additionally, we introduce a new back-verification technique that uses LLMs to validate the correctness of the generated answers. Furthermore, we employ a pruning tree search to optimize search time while achieving strong performance. Notably, our method improves Qwen2-7b-Instruct’s score from 36.94 to 61.52 (outperforming GPT-4’s 42.5) on the MATH benchmark. The code is made available at https://github.com/Aurora-slz/BEATS
@@ 1 Introduction
LLMs have demonstrated exceptional performance across diverse tasks and domains Touvron et al. (2023); meta llama (2024); Bai et al. (2023a), excelling in zero-shot and few-shot scenarios. Recent advancements in scaling laws and fine-tuning have further enhanced their capabilities, enabling their application in complex real-world tasks such as natural language understanding and multimodal processing.
Among the various capabilities of LLMs, mathematical proficiency is crucial, as it reflects not only logical reasoning but also the model’s capacity for structured problem-solving. Mastery of mathematical tasks necessitates precision, adherence to complex rules, and the application of algorithms, all of which are essential indicators of an LLM’s overall reasoning and cognitive abilities. There are generally two approaches to enhance mathematical capability. The first set of methods trains LLMs to improve their mathematical skills. Models such as Mammoth Yue et al. (2023; 2024) and internLM math, along with DeepSeek Shao et al. (2024), utilize vast amounts of data to develop robust mathematical models. The second set of methods employs tree search and self-correction techniques to enhance mathematical abilities. Techniques like ToT Yao et al. (2024), RAP Hao et al. (2023), ReST-MCTS* Zhang et al. (2024), and LiteSearch Wang et al. (2024) leverage tree structures and search methods such as DFS and MCTS. However, both approaches still encounter suboptimal results. They face the following challenges:
@@
Self-improving models typically split tasks into multiple subtasks for LLMs to solve. However, previous works such as  Yao et al. (2024); Wang et al. (2024) often neglect the design of curated prompts, particularly those that include rewriting questions to eliminate ambiguities.
@@
Previous works like  Yao et al. (2024); Wang et al. (2024) typically employ voting-based verification methods. However, they overlook the fact that LLMs can make the same mistakes across multiple routes.
High Computational Cost
Previous works utilizing pre-training or SFT techniques  Yue et al. (2023; 2024); Ying et al. (2024) often suffer from insufficient amounts of data and high computational costs. Search-based methods like  Yao et al. (2024); Wang et al. (2024), while introducing MCTS to reduce computational costs, still encounter challenges due to the necessity of deep trees to achieve optimal performance.
@@
To address these challenges, we propose BEATS, a novel method for efficient search aimed at enhancing mathematical performance. Our method guides the model to answer problems step by step, thereby avoiding ambiguities in problem statements. We meticulously design prompts that instruct the model to rewrite, solve one step at a time, and generate answers based on preceding steps. Additionally, traditional verification methods in tree search, such as majority voting, may be unreliable, as LLMs can perpetuate the same mistakes across multiple branches. To overcome this, we introduce a back-verification technique that re-submits both the answer and the problem to the model for a judgment of correctness, leveraging the model’s capabilities while reducing its reasoning difficulty. Furthermore, we employ a pruning tree search to optimize search time while achieving strong performance. It is worth noting that with our meticulously designed pruning tree, we can control search time; simultaneously, compared to Monte Carlo Tree Search (MCTS), the pruning tree is able to search through every leaf node, ensuring performance, while MCTS is more likely to search based on prior experience.
@@
  • Meticulously Designed Prompt We have developed three new curated prompts to solve math problems step by step, thereby avoiding ambiguities in the problem statements.
  • Pruning Tree Search for Controllable Inference Time We implement a pruning strategy for the tree by imposing constraints on the search steps. Specifically, we restrict the rewriting of the question to once and the provision of the answer to once.
  • New Effective Verification Method We propose a novel back-verification method that re-submits both the answer and the problem to the model for a judgment of correctness, as shown in Figure 1. This approach enhances the performance of searching in LLMs compared to majority voting.
@@ 2.1 Math Large Language Models
Large Language Models (LLMs) have demonstrated significant capabilities across various tasks, including mathematical problem-solving, which is a critical skill for these models. However, learning to solve mathematical problems poses challenges for LLMs, often requiring large amounts of training data and substantial computational resources. In this paper, we review several state-of-the-art (SOTA) models specifically designed to tackle mathematical problems.
Llemma Azerbayev et al. (2021) integrates both code and mathematical data to train models, resulting in strong performance. InternLM2 Ying et al. (2024) utilizes a vast amount of math-related pre-training corpus to achieve high performance. Mammoth Yue et al. (2023) collected Chain-of-Thought (CoT) data for fine-tuning language models and achieved impressive results. Mammoth2 Yue et al. (2024) builds on Mammoth by collecting WebInstruct, one of the largest open-source math datasets, and uses it to fine-tune LLMs, resulting in state-of-the-art (SOTA) performance. DeepSeek Shao et al. (2024) employs preference-based mathematical data to perform an additional stage of reinforcement learning, achieving SOTA results.
In addition to models explicitly trained for mathematics, a few foundation models exhibit exceptional mathematical proficiency. Llama3 Touvron et al. (2023) has shown remarkable performance in solving mathematical problems. Qwen2 Bai et al. (2023b), another series of outstanding models, is one of the SOTA open-source models. Furthermore, closed-source models like Claude and GPT also demonstrate strong capabilities in mathematical problem solving.
@@ 2.2 Prompt Engineering for Large Language Models
The effectiveness of large language models in various applications largely depends on the quality of the prompts used. There are already many designed prompts that can significantly enhance the performance of LLMs Kojima et al. (2022); Wei et al. (2022); Yao et al. (2024); Besta et al. (2024); Yang et al. (2024); Wang et al. (2023). However, these methods that rely on manual prompt engineering are far less scalable. In the field of mathematical logical reasoning for LLMs, the Chain of Thought and its derived strategies are widely popular due to their effectiveness. Zero-shot CoT Kojima et al. (2022) is adding a simple sentence like “Let’s think step by step” at the end of questions to assist LLMs in generating reasoning steps. Instead of Zero-shot CoT, Manual-Cot Wei et al. (2022) provides reasoning steps as few shots. Self-Consistency further improves language models’ reasoning performance by generating a diverse set of reasoning paths and choosing the most consistent answer in the final answer set. Tree of Thought Yao et al. (2024) and GOT Besta et al. (2024) extend the reasoning pathway from linear to non-linear data structures by leveraging multiple LLM queries to elicit different plausible reasoning paths Yang et al. (2024). Buffer of Thought (BOT) Yang et al. (2024) designs a series of thought-template for tasks, and for each problem, it retrieve a relevant thought-template to prompt LLMs. PS prompting Wang et al. (2023) improves COT by encouraging LLMs to devise a plan before attempting to solve a problem.
@@ 2.3 Reasoning in Large Language Models
The recently introduced GPT-o1 has shown exceptional results in solving mathematical problems, largely due to the integration of a novel reasoning module. Our tree search method can be considered a mathematical reasoning approach. In this paper, we present a comprehensive review of existing reasoning methods for LLMs. Li et al. Li et al. (2024) demonstrated that, with Chain of Thought (COT), LLMs can achieve arbitrarily strong performance. Zelikman et al. Zelikman et al. (2024) also revealed that it is possible for LLMs to ”think before reasoning.” To facilitate such thinking, tree structures and verification mechanisms are typically employed. Tree of Thought Yao et al. (2024) leverages tree search and majority voting to enhance inference performance. Building upon this work, Zhang et al. Zhang et al. (2024) employed Monte Carlo Tree Search to achieve efficient and effective tree search.
Several other works fine-tune LLMs to enable self-improvement capabilities. For instance, Chen et al. Chen et al. (2024b) employed Step-Level Value Preference Optimization to achieve strong model performance. In a related work, Chen et al. Chen et al. (2024a) trained value and policy functions and applied step-level beam search during inference to enhance mathematical abilities. Kumar et al. Kumar et al. (2024) utilized reinforcement learning and oracle feedback to teach models self-correction.
@@ 3 Method
Figure 2: Visualization of our BEATS method.
@@ 3.1 Prompt Design
The prompt is illustrated in Figure 5(c). This prompt emphasizes reformulating the initial query to enhance clarity and specificity, thereby facilitating a more effective search process. This approach is novel, as many problem descriptions are frequently ambiguous or unclear, leading to incorrect answers. For example, the query, Josh decides to try flipping a house. He buys a house for \$80,000 and then invests \$50,000 in repairs. This increased the value of the house by 150%. How much profit did he make?, can introduce ambiguity. By incorporating a step to rewrite questions, we aim to eliminate such ambiguities, ensuring that the model fully comprehends the problem before attempting to solve it. This helps prevent errors that result from misinterpretations of the initial query.
@@ 3.2 Pruning Tree Search
As shown in Figure 2, a node in the tree is denoted by $u_{d}$, where $d$ indicates the depth of the node. For a given node $u_{d}$, its ancestor nodes up to the root are denoted by the sequence $u_{d-1},...,u_{1}$. Each node is associated with a prompt that concatenates the responses from previous rounds. These prompts, containing prior rounds of answers, are fed into the action module to generate further responses leading to the correct answer.
@@ 4 Experiment
Zero-Shot CoT LLaMA3 8B 27.8 50.27
CoT Yi-1.5 6B 30.42 64.47
CoT Qwen2 7B 36.94 76.63
Hard Voting@8 LLaMA3 8B 30.00 78.39
Hard Voting@64 LLaMA3 8B 33.00 83.24
SFT WizardMath LLaMA2 7B 10.7 54.9
MuggleMath LLaMA2 7B - 68.40
MetaMath LLaMA2 7B 19.80 66.50
LEMA-LLaMA LLaMA2 7B 9.40 54.10
Search ToT LLaMA3 8B 13.60 69.07
RAP LLaMA3 8B 18.80 80.59
@@
LiteSearch LLaMA3 8B - 82.30
Llama-2+M* (BS@16) LLaMA2 13B 32.40 66.30
@@
Search Ours (w.o. BackVerify) LLaMA3 8B 35.17 83.62
Ours LLaMA3 8B 42.93 88.48
Ours (w.o. BackVerify) Yi-1.5 6B 42.01 74.68
Ours Yi-1.5 6B 51.27 76.12
Ours (w.o. BackVerify) Qwen2 7B 57.28 81.50
Ours Qwen2 7B 61.52 83.02
@@
Zero-Shot CoT LLaMA3 8B 53.90 21.20 27.35
CoT Yi-1.5 6B 76.40 34.63 38.39
CoT Qwen2 7B 85.2 32.68 53.36
SFT Code-Llama - 13B 60.00 3.80 27.60
WizardMath LLaMA2 13B 51.90 14.90 36.10
Platypus LLaMA2 13B 55.40 7.40 42.30
Platypus LLaMA1 30B+ 51.70 13.60 40.50
Platypus LLaMA2 65B+ 51.80 21.70 48.10
Ocra-Platypus LLaMA2 13B 56.80 7.90 35.30
MAmmoTH LLaMA2 13B 72.40 43.20 61.20
MAmmoTH-Coder Code-Llama 13B 73.70 47.10 66.40
Galactica GAL 30B 41.60 13.20 34.70
Tulu LLaMA2 30B+ 59.00 10.30 43.40
Guanaco LLaMA2 65B+ 66.80 20.20 40.50
Search Ours (w.o. BackVerify) LLaMA3 8B 80.60 72.76 66.99
Ours LLaMA3 8B 88.70 78.40 73.61
Ours (w.o. BackVerify) Yi-1.5 6B 79.30 34.72 75.43
Ours Yi-1.5 6B 83.70 34.82 77.93
Ours (w.o. BackVerify) Qwen2 7B 88.80 35.21 72.84
Ours Qwen2 7B 90.70 36.19 73.16

### v2

@@ Abstract
Large Language Models (LLMs) have exhibited exceptional performance across a broad range of tasks and domains. However, they still encounter difficulties in solving mathematical problems due to the rigorous and logical nature of mathematics. Previous studies have employed techniques such as supervised fine-tuning (SFT), prompt engineering, and search-based methods to improve the mathematical problem-solving abilities of LLMs. Despite these efforts, their performance remains suboptimal and demands substantial computational resources. To address this issue, we propose a novel approach, BEATS, to enhance mathematical problem-solving abilities. Our method leverages newly designed prompts that guide the model to iteratively rewrite, advance by one step, and generate answers based on previous steps. Additionally, we employ a pruning tree search to optimize search time while achieving strong performance. Furthermore, we introduce a new back-verification technique that uses LLMs to validate the correctness of the generated answers. Notably, our method improves Qwen2-7b-Instruct’s score from 36.94 to 61.52 (outperforming GPT-4’s 42.5) on the MATH benchmark. The code is made available at https://github.com/Aurora-slz/BEATS
@@ 1 Introduction
LLMs have demonstrated exceptional performance across diverse tasks and domains (Touvron et al., 2023; meta llama, 2024; Bai et al., 2023a), excelling in zero-shot and few-shot scenarios. Recent advancements in scaling laws and fine-tuning have further enhanced their capabilities, enabling their application in complex real-world tasks such as natural language understanding and multimodal processing.
Among the various capabilities of LLMs, mathematical proficiency is crucial, as it reflects not only logical reasoning but also the model’s capacity for structured problem-solving. Mastery of mathematical tasks necessitates precision, adherence to complex rules, and the application of algorithms, all of which are essential indicators of an LLM’s overall reasoning and cognitive abilities. There are generally two approaches to enhance mathematical capability. The first set of methods trains LLMs to improve their mathematical skills. Models such as Mammoth (Yue et al., 2023; 2024) and Internlm-math (Ying et al., 2024), along with DeepSeek (Shao et al., 2024), utilize vast amounts of data to develop robust mathematical models. The second set of methods employs tree search and self-correction techniques to enhance mathematical abilities. Techniques like ToT (Yao et al., 2024), RAP (Hao et al., 2023), ReST-MCTS* (Zhang et al., 2024), and LiteSearch (Wang et al., 2024) leverage tree structures and search methods such as BFS, DFS and Monte Carlo Tree Search (MCTS). However, both approaches still encounter suboptimal results. They face the following challenges:
@@
Self-improving models (Yao et al., 2024; Wang et al., 2024) typically address problems by either decomposing them into subproblems or rewriting them, followed by solving through methods CoT or Process of Thought (PoT). However, they tend to overlook the issue of ambiguous problem statements. As illustrated by the root node in Figure 1(a), vague expressions can mislead the LLM’s understanding.
High Computational Cost
Previous researches utilizing pre-training or SFT techniques  (Yue et al., 2023; 2024; Ying et al., 2024) often suffer from insufficient amounts of data and high computational costs. Search-based approaches enhance mathematical reasoning during the inference stage, thus avoiding the pressure of additional training. However, due to the vast search space, a naive search algorithm can lead to a significant increase in inference time(Yao et al., 2024). Although Wang et al. (2024) employs MCTS to compress the search space, which may result in the absent of correct answers.
@@
When selecting among multiple candidate answers to a problem, previous works like  Yao et al. (2024); Wang et al. (2024) typically employ voting-based verification methods. However, they overlook the fact that LLMs can make the same mistakes across multiple routes.
@@
To address these challenges, we propose BEATS, a novel method for efficient search aimed at enhancing mathematical performance. Our method guides the model to answer problems instructed by clarified question, thereby avoiding ambiguities in problem statements. We meticulously design prompts that instruct the model to disambiguate, solve one step at a time, and directly generate answers based on preceding steps. Additionally, traditional verification methods in tree search, such as majority voting, may be unreliable, as LLMs can perpetuate the same mistakes across multiple branches. To overcome this, we introduce a back-verification technique that re-submits both the answer and the problem to the model for a judgment of correctness, leveraging the model’s capabilities while reducing its reasoning difficulty. Furthermore, we employ a pruning tree search to optimize search time while achieving strong performance. It is worth noting that with our meticulously designed pruning tree, we can control search expenses; simultaneously, compared to MCTS, the pruning tree is able to search through every leaf node, ensuring promising performance, while MCTS is more likely to search based on prior experience.
@@
  • Meticulously Designed Prompt We developed three newly curated prompts designed to solve mathematical problems step-by-step, provide final answers, and, most importantly, avoiding ambiguities in problem statements.
  • Pruning Tree Search for Controllable Inference Time We implement a pruning strategy for the tree by imposing constraints on the search steps. Specifically, we restrict the rewriting of the question to once and terminate the tree construction when answer is achieved.
  • New Effective Verification Method We propose a new back-verification method that re-submits both the answer and the problem to the model for a judgment of correctness, as shown in Figure 1. This approach enhances the performance of searching in LLMs compared to majority voting.
@@ 2.1 Math Large Language Models
LLMs have demonstrated significant capabilities across various tasks, including mathematical problem-solving, which is a critical skill for these models. However, learning to solve mathematical problems poses challenges for LLMs, often requiring large amounts of training data and substantial computational resources. In this paper, we review several state-of-the-art (SOTA) models specifically designed to tackle mathematical problems.
Llemma (Azerbayev et al., 2021) integrates both code and mathematical data to train models, resulting in strong performance. InternLM2 (Ying et al., 2024) utilizes a vast amount of math-related pre-training corpus to achieve high performance. Mammoth (Yue et al., 2023) collected Chain-of-Thought (CoT) data for fine-tuning language models and achieved impressive results. Mammoth2 (Yue et al., 2024) builds on Mammoth by collecting WebInstruct, one of the largest open-source math datasets, and uses it to fine-tune LLMs, resulting in SOTA performance. DeepSeek (Shao et al., 2024) employs preference-based mathematical data to perform an additional stage of reinforcement learning, achieving SOTA results.
In addition to models explicitly trained for mathematics, a few foundation models exhibit exceptional mathematical proficiency. Llama3 (Touvron et al., 2023) has shown remarkable performance in solving mathematical problems. Qwen2 (Bai et al., 2023b), another series of outstanding models, is one of the SOTA open-source models. Furthermore, closed-source models like Claude and GPT also demonstrate strong capabilities in mathematical problem solving.
@@ 2.2 Prompt Engineering for Large Language Models
The effectiveness of large language models in various applications largely depends on the quality of the prompts used. There are already many designed prompts that can significantly enhance the performance of LLMs (Kojima et al., 2022; Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Yang et al., 2024; Wang et al., 2023a). However, these methods that rely on manual prompt engineering are far less scalable. In the field of mathematical logical reasoning for LLMs, the Chain of Thought and its derived strategies are widely popular due to their effectiveness. Zero-shot CoT (Kojima et al., 2022) is adding a simple sentence like “Let’s think step by step” at the end of questions to assist LLMs in generating reasoning steps. Instead of Zero-shot CoT, Manual-Cot (Wei et al., 2022) provides reasoning steps as few shots. Self-Consistency further improves language models’ reasoning performance by generating a diverse set of reasoning paths and choosing the most consistent answer in the final answer set. Tree of Thought (Yao et al., 2024) and GOT (Besta et al., 2024) extend the reasoning pathway from linear to non-linear data structures by leveraging multiple LLM queries to elicit different plausible reasoning paths (Yang et al., 2024). Buffer of Thought (BOT) (Yang et al., 2024) designs a series of thought-template for tasks, and for each problem, it retrieve a relevant thought-template to prompt LLMs. PS prompting (Wang et al., 2023a) improves COT by encouraging LLMs to devise a plan before attempting to solve a problem. In this paper, we employ meticulously designed prompts to enhance the model’s mathematical capabilities.
@@ 2.3 Reasoning in Large Language Models
The recently introduced GPT-o1 model has demonstrated outstanding performance in solving mathematical problems, primarily due to its integration of a novel reasoning module. Our proposed tree search methodology can be categorized as a mathematical reasoning technique. In this paper, we provide a comprehensive review of existing reasoning methods for LLMs. Li et al. Li et al. (2024c) showed that LLMs can achieve arbitrarily high performance with Chain-of-Thought (CoT) prompting. Similarly, Zelikman et al. Zelikman et al. (2024) highlighted the potential for LLMs to “think before reasoning,” facilitated by the use of tree structures and verification mechanisms. Tree of Thought (Yao et al., 2024) leverages tree search and majority voting to improve inference performance. Building on this foundation, Zhang et al. Zhang et al. (2024) applied Monte Carlo Tree Search (MCTS) to achieve efficient and effective tree-based search.
Other works have focused on fine-tuning LLMs to develop self-improvement capabilities. For example, Chen et al. Chen et al. (2024b) employed Step-Level Value Preference Optimization to achieve high model performance. Another related work, AlphaMath Chen et al. (2024a), proposed by Chen et al., utilized value and policy functions along with step-level beam search during inference to enhance mathematical problem-solving abilities. Kumar et al. Kumar et al. (2024) further employed reinforcement learning and oracle feedback to train models for self-correction.
@@ 3 Method
Figure 2: Visualization of the search algorithm in BEATS.
@@ 3.1 Prompt Design
The prompt is illustrated in Figure 7(c). This prompt emphasizes reformulating the initial query to enhance clarity and specificity, thereby facilitating a more effective search process. This approach is necessary, as many problem descriptions are frequently ambiguous or unclear, leading to incorrect answers. For example, the query, Josh decides to try flipping a house. He buys a house for \$80,000 and then invests \$50,000 in repairs. This increased the value of the house by 150%. How much profit did he make?, can introduce ambiguity. By incorporating a step to rewrite questions, we aim to eliminate such ambiguities, ensuring that the model fully comprehends the problem before attempting to solve it. This helps prevent errors that result from misinterpretations of the initial query.
@@ 3.2 Pruning Tree Search
As shown in Figure 2 and Algorithm 1, a node in the tree is denoted by $u_{d}$, where $d$ indicates the depth of the node. For a given node $u_{d}$, its ancestor nodes up to the root are denoted by the sequence $u_{d-1},...,u_{1}$. Each node is associated with a prompt that concatenates the responses from previous rounds. These prompts, containing prior rounds of answers, are fed into the action module to generate further responses leading to the correct answer.
@@ 4 Experiment
Zero-Shot Chain-of-Thought LLaMA3 8B 27.80 50.27
Chain-of-Thought Yi-1.5 6B 30.42 64.47
Chain-of-Thought Qwen2 7B 36.94 76.63
Hard Voting@8 (Wang et al., 2024) LLaMA3 8B 30.00 78.39
Hard Voting@64 (Wang et al., 2024) LLaMA3 8B 33.00 83.24
SFT WizardMath (Luo et al., 2023) LLaMA2 7B 10.70 54.90
MuggleMath (Li et al., 2024b) LLaMA2 7B - 68.40
MetaMath (Yu et al., 2023) LLaMA2 7B 19.80 66.50
LEMA-LLaMA (An et al., 2023) LLaMA2 7B 9.40 54.10
Search ToT (Yao et al., 2024) LLaMA3 8B 13.60 69.07
RAP (Hao et al., 2023) LLaMA3 8B 18.80 80.59
@@
LiteSearch (Wang et al., 2024) LLaMA3 8B - 82.30
Llama-2+M* (BS@16) (Kang et al., 2024) LLaMA2 13B 32.40 66.30
@@
Search BEATS (w.o. BackVerify) LLaMA3 8B 35.17 83.62
BEATS LLaMA3 8B 42.93 88.48
BEATS (w.o. BackVerify) Yi-1.5 6B 42.01 74.68
BEATS Yi-1.5 6B 51.27 76.12
BEATS (w.o. BackVerify) Qwen2 7B 57.28 81.50
BEATS Qwen2 7B 61.52 83.02
@@
Zero-Shot Chain-of-Thought LLaMA3 8B 53.90 21.20 27.35
Chain-of-Thought Yi-1.5 6B 76.40 34.63 38.39
Chain-of-Thought Qwen2 7B 85.20 32.68 53.36
SFT Code-Llama (Roziere et al., 2023) - 13B 60.00 3.80 27.60
WizardMath (Luo et al., 2023) LLaMA2 13B 51.90 14.90 36.10
Platypus (Lee et al., 2023) LLaMA2 13B 55.40 7.40 42.30
Platypus (Lee et al., 2023) LLaMA1 30B+ 51.70 13.60 40.50
Platypus (Lee et al., 2023) LLaMA2 65B+ 51.80 21.70 48.10
Ocra-Platypus (Lee et al., 2023) LLaMA2 13B 56.80 7.90 35.30
MAmmoTH (Yue et al., 2023) LLaMA2 13B 72.40 43.20 61.20
MAmmoTH-Coder (Yue et al., 2023) Code-Llama 13B 73.70 47.10 66.40
Galactica (Taylor et al., 2022) GAL 30B 41.60 13.20 34.70
Tulu (Wang et al., 2023b) LLaMA2 30B+ 59.00 10.30 43.40
Guanaco (Dettmers et al., 2023) LLaMA2 65B+ 66.80 20.20 40.50
Search BEATS (w.o. BackVerify) LLaMA3 8B 80.60 72.76 66.99
BEATS LLaMA3 8B 88.70 78.40 73.61
BEATS (w.o. BackVerify) Yi-1.5 6B 79.30 34.72 75.43
BEATS Yi-1.5 6B 83.70 34.82 77.93
BEATS (w.o. BackVerify) Qwen2 7B 88.80 35.21 72.84
BEATS Qwen2 7B 90.70 36.19 73.16